As a restaurant owner there are many decisions that you need to make on a daily basis, such as where to keep inventory levels. You wish to replenish your stock of dishes by purchasing 250 sets for your restaurant. You have two dish design from which to choose. One design costs $20 per set and the other $45 per set. If you only have $6,800 to spend, how many of each design should you order? I need help setting this up.
Answer by stanbon(75887)

You wish to replenish your stock of dishes by purchasing 250 sets for your restaurant. You have two dish design from which to choose. One design costs $20 per set and the other $45 per set. If you only have $6,800 to spend, how many of each design should you order?
Equations:
a + b = 250
20a + 45b = 6800
----
Multiply thru 1st by 20 to get:
20a + 20b = 20*250
Subtract that from the 2nd eq.
25b = 1800
b = 72 (# of sets of $20 dishes to purchase)
a = 250-72 = 178 (# of sets of $45 dishes to purchase)
===================
Cheers,
Stan H.